# Mind the Prompt: A Novel Benchmark for Prompt-based Class-Agnostic Counting

### v1

@@ Abstract
Class-agnostic counting (CAC) is a recent task in computer vision that aims to estimate the number of instances of arbitrary object classes never seen during model training. With the recent advancement of robust vision-and-language foundation models, there is a growing interest in prompt-based CAC, where object categories to be counted can be specified using natural language. However, we identify significant limitations in current benchmarks for evaluating this task, which hinder both accurate assessment and the development of more effective solutions. Specifically, we argue that the current evaluation protocols do not measure the ability of the model to understand which object has to be counted. This is due to two main factors: (i) the shortcomings of CAC datasets, which primarily consist of images containing objects from a single class, and (ii) the limitations of current counting performance evaluators, which are based on traditional class-specific counting and focus solely on counting errors. To fill this gap, we introduce the Prompt-Aware Counting (PrACo) benchmark, which comprises two targeted tests, each accompanied by appropriate evaluation metrics. We evaluate state-of-the-art methods and demonstrate that, although some achieve impressive results on standard class-specific counting metrics, they exhibit a significant deficiency in understanding the input prompt, indicating the need for more careful training procedures or revised designs. The code for reproducing our results is available at https://github.com/ciampluca/PrACo.
@@ 1 Introduction
Object counting is the automated process of estimating the number of objects in still images or video frames [14]. Very recently, research about this task has undergone a detour towards class-agnostic counting (CAC) aiming at enumerating the instances of arbitrary object categories never seen during model training [21]. This new trend overcomes the limitations affecting class-specific counting methods that instead estimate the number of objects belonging to specific classes, requiring individually trained networks for each object type such as people [22, 24, 18, 4, 10], vehicles [29, 1, 6], or cells [8, 5]. In contrast, class-agnostic counting draws inspiration from humans’ instinctive ability to discern what merits counting when confronted with unfamiliar objects, and it enables users to specify arbitrary categories of interest at inference time without needing model retraining and gathering new annotated data.
Figure 1: Prompt-guided counting models – CounTX [2] in this example – have problems in correctly understanding the user-provided text containing the class of objects to be counted. The confusion affects classes very well semantically separated – like marbles and elephants. The plot shows that classes exist that are not present in the image, for which state-of-the-art models estimate a higher count than the ground-truth class (in orange).
@@
Target object classes in CAC can be specified with visual exemplars, i.e., bounding boxes surrounding object samples in the image [25, 27, 17, 9, 28], or with text prompts, i.e., textual object descriptions [26, 2, 11, 12, 23, 19, 16]. Even if current techniques exploiting visual exemplars outperform prompt-based methods, the latter setting is very appealing since it leverages the model’s natural language understanding, enhancing flexibility as less human intervention in the loop is required, i.e., humans do not have to specify bounding box exemplars as input [2, 19].
@@
However, in this paper, we figured out that current benchmarks for assessing prompt-based CAC approaches are affected by severe limitations hampering their proper evaluation and the development of new, more performing solutions. As shown in Fig. 1, we empirically verified that some state-of-the-art prompt-based CAC approaches are not always able to truly understand which object class has to be counted from the textual description, even if they reach top results on current performance evaluators. In contrast, they tend to count object instances belonging to the predominant class despite the prompt meaning. This is a critical failure in real-world scenarios if we imagine that such systems are usually deployed to count specific objects – e.g., imagine a surveillance scenario in which the officer wants to count the number of pedestrians transiting on the street, but at that moment, only cars are present. We argue that this lack relies on two key factors: (i) the shortcomings affecting the main CAC datasets and (ii) the limitations affecting current counting performance evaluators. Specifically, most of the samples making up the datasets suitable for CAC contain objects belonging to a single class, making it hard to evaluate the model’s ability to discriminate different object classes within an image. Furthermore, the metrics exploited for measuring the performance of CAC models are inherited from class-specific counting and do not take into account some key factors of prompt-based CAC, focusing only on the counting error and disregarding to assess the model’s ability to understand the object category described by the textual prompt.
@@
In this work, we propose Prompt-Aware Counting (PrACo), a novel benchmark designed to quantitatively evaluate the effectiveness of prompt-based CAC approaches in accurately interpreting input prompts. The benchmark includes two specific tests tackling the two above-mentioned limitations: (i) a negative-label test, which probes single-class images by using prompts that refer to absent classes, and (ii) a mosaic test, which evaluates artificially created images containing two distinct classes, with the prompt describing only one of them. Furthermore, PrACo comes with ad-hoc evaluation metrics specific for each of these two tests, overcoming the limitations affecting class-specific counting performance evaluators. To validate our new benchmark, we assess several recent SOTA prompt-based CAC techniques, and we show that some of them show a notable weakness in understanding the object class textual descriptions despite achieving top performance on standard class-specific metrics. This highlights the necessity for more refined training procedures or reconsideration of their architectural designs.
@@
  • We propose PrACo, a novel benchmark aiming at assessing prompt-based CAC methods to overcome the limitations affecting current datasets and metrics hampering the development of more effective solutions.
  • Contextually to this benchmark, we introduce a set of ad-hoc evaluation protocols and associated metrics to evaluate the models’ ability to count object classes described with natural language.
  • We perform an experimental evaluation by exploiting PrACo showing that many recent state-of-the-art methods exhibit a remarkable deficiency in understanding the object to be counted although obtaining top results in standard class-specific counting metrics.
@@ 2.1 Class-specific Object Counting
Class-specific object counting is one of the core tasks in computer vision due to its widespread applicability to many real-world applications. Thus, many methods have been proposed to count specific object categories, such as people [22, 24, 18, 4, 10], cars [29, 1, 6], insects [3, 7], or cells [8, 5]. Approaches that regress and sum density maps have emerged to be more accurate in crowded scenarios [22, 24, 18, 4, 10, 3] than techniques relying on prior detection of object instances [1, 6]. Either way, the main drawback of class-specific counting lies in requiring individually trained networks and, consequently, labeled datasets for each object type.
@@ 2.2 Prompt-based Class-agnostic Counting
Very recently, class-agnostic counting (CAC) has generalized class-specific counting to open-world settings, where users can specify arbitrary object classes never seen during model training [21]. Specifically, the object category to be counted is selected at inference time by exploiting (i) visual exemplars, i.e., some bounding boxes around the objects of interest within the same input image [25, 27, 17, 9, 28], or (ii) textual prompts, i.e., natural-language descriptions of the object class [26, 2, 11, 12, 23, 19, 16]. The first setting provides the best performance: methods such as CACViT [25], SSD [27], CounTR [17], LOCA [9], and SAFECount [28] reach state-of-the-art in all CAC benchmarks. In contrast, prompt-based CAC approaches allow the input of object properties through text, and they are particularly appealing since they enhance overall flexibility.
@@
Most of the state-of-the-art prompt-based CAC approaches rely on vision-language foundation models that map text-image pairs to a joint embedding space; the produced features are then passed to a decoder in charge of producing density maps. The authors in [26] trained a conditional variational autoencoder (VAE) based on the popular CLIP model [20] to generate visual exemplar prototypes conditioned with their semantic embedding encoded in the provided category name; prototypes are then used with existing visual exemplar-based CAC techniques. Concurrently, [2] proposed CounTX that instead is trained end-to-end and produces object counts directly, skipping the intermediate visual exemplar proposal step; furthermore, it accepts a more detailed specification of the target object to count (rather than simply using a class name). More in detail, it is built on top of a CLIP model, where a feature interaction module is used to mine the similarities between image patches and class descriptions. Similarly, CLIP-Count [11] relies on CLIP to align the text embedding with dense visual features. However, in this paper, the authors also designed a hierarchical patch-text interaction module to propagate semantic information across different resolution levels of visual features. Another similar CLIP-based model trained end-to-end is represented by VLCounter [12]. Here, the authors incorporated three modules to efficiently finetune CLIP for the counting task and to exploit intermediate features across different encoding layers of CLIP in the decoding stage. Very recently, DAVE [19] proposed a two-stage detect-and-verify paradigm. In the first stage, they estimate a high-recall set of candidate detections, which are then analyzed and filtered in the second verification step, relying on unsupervised clustering and CLIP embedding. Differently from the above-described architectures, TFPOC [23] employed a detection-based technique, leveraging the popular SAM [13] for instance segmentation. The authors proposed a two-stage approach. In the first stage, they exploited CLIP-Surgery [15], an enhanced version of CLIP, to produce visual exemplar prototypes by computing the similarity between images and text representations. In the second step, they computed a similarity map by using cosine similarity between image features and the masks corresponding to the reference objects produced by SAM prompted with the bounding boxes from the first stage. However, in this work, we empirically demonstrate that these prompt-based CAC models are not always able to truly understand the object category to be counted from the textual description, tending simply to the most predominant object class instead.
@@ 2.3 Dataset and Metrics
The gold standard dataset for CAC is FSC-147 [21]. It comprises 6,135 images covering 147 object categories. Annotations consist of dots over the approximate centroids of each object instance, as usual for the counting task – ground truth density maps are created by superimposing Gaussian kernels centered on these dots. Furthermore, the authors provided three bounding boxes per image that localize the exemplars and the natural language name of the object category to which they belong – the category is unique for each image, and it has been chosen arbitrarily if multiple categories were present. However, cases in which multiple object categories are present in the same image are just a few, and this represents an inherent limitation of FSC-147, making it hard to evaluate the model’s ability to discriminate different object classes within an image.
Standard counting metrics in the literature are the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE), defined as $\text{MAE}=\frac{1}{N}\sum_{n=1}^{N}\left|\tilde{c}_{n}-c_{n}\right|$ and $\text{RMSE}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(\tilde{c}_{n}-c_{n})^{2}}$, where $N$ is the number of test images, and $\tilde{c}_{n}$ and $c_{n}$ are the ground truth and predicted counts, respectively. However, these evaluators were inherited from the class-specific counting task and exhibit severe limitations particularly evident in prompt-based CAC. Specifically, they do not take into account which object category has to be counted nor the model’s ability to understand the provided textual prompt.
@@ 3.1 Overview
Standard class-specific counting benchmarks have several limitations when used to evaluate prompt-based CAC models. In this context, it is essential to not only measure the model’s accuracy in estimating the number of object instances but also to verify whether the model correctly identifies which object should be counted based on the given textual prompt.
We assume to have a collection of $N$ tuples $\mathcal{X}=\{(I_{1},P_{1}),(I_{2},P_{2}),\dots,(I_{N},P_{N})\}$, where $I_{i}$ is an image and $P_{i}$ is the class name of the objects within $I_{i}$, i.e. each image contains objects belonging to only a single class. We define $P_{i}$ as the positive class and all the $\{P_{j}\}_{j\neq i}$ as the negative classes for the given $I_{i}$. Without loss of generality and for simplifying the formalisms, we are assuming that each class appears in only one image of the dataset, although in real-world scenarios, datasets collect multiple images for the same class of objects.
On this common structure, PrACo introduces two test suites, which are detailed in the following sections: (i) a negative-label test and (ii) a mosaic test.
@@ 3.2 Negative-label test
The negative-label test evaluates single-class images by prompting the model with textual descriptions referring to absent object classes. Formally, we are given a counting model $\mathcal{M}(I,P)$, which takes as input the image $I$ and an arbitrary prompt $P$ containing the class to be counted and outputs an estimated count for that class. For all the elements $\{I_{i},P_{j}\}_{i,j=1}^{N}$, we compute $c_{ij}=\mathcal{M}(I_{i},P_{j})$ to obtain the count that the model predicts for each image when prompted with all the available classes in the dataset. In order for this test to perfectly succeed, we would ideally require to achieve the following $c_{ij}$ configuration:
@@
where $\tilde{c}_{i}$ is the ground truth count for the image $I_{i}$. In other words, other than solely checking that the model correctly estimates the number of objects for the positive class – which is what classical counting metrics check – we are also required to verify that the model estimates a near-zero count when prompted with negative classes. The ideal situation is illustrated in the left part of Figure 2. To quantitatively assess this test, we introduce two metrics that, unlike the classical counting metrics, try to quantify the error caused by counting the wrong objects.
@@
NMN is the absolute counting error computed by prompting the model with the negative classes, normalized by the ground truth of the positive class. It is computed as follows (more details in the supplementary):
@@
PCCN is an aggregate measure defined as the percentage of samples in the dataset where the model produces a positive count estimate that is nearer to the ground truth with respect to the mean of the negative classes. Formally, it is defined as follows:
@@
where $d^{\text{neg}}_{i}=|\frac{1}{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}c_{ij}-\tilde{c}_{i}|$ is the absolute difference between the mean of the model’s predictions based on the negative labels and the ground truth count for the $i$-th image, $d^{\text{pos}}_{i}=|c_{ii}-\tilde{c_{i}}|$ is the absolute difference between the model’s prediction based on the positive label and the ground truth count for the $i$-th image, and $\mathbb{I}(\cdot)$ is the indicator function, which equals 1 if the condition inside it is true, and 0 otherwise. Unlike NMN, this metric gives an interesting global view by mixing positive and negative label predictions. It provides an overall quantitative assessment of strong failures in the model. In fact, a mean negative estimate closer to ground truth than the positive one may signify that the model is strongly biased to count most of the negative classes instead of the positive one, which in turn may be due to the model completely losing the semantics of the given prompt. The higher the WPR, the better the resulting model.

### v2

@@ Abstract
Recently, object counting has shifted towards class-agnostic counting (CAC), which counts instances of arbitrary object classes never seen during model training. With advancements in robust vision-and-language foundation models, there is a growing interest in prompt-based CAC, where object categories are specified using natural language. However, we identify significant limitations in current benchmarks for evaluating this task, which hinder both accurate assessment and the development of more effective solutions. Specifically, we argue that the current evaluation protocols do not measure the ability of the model to understand which object has to be counted. This is due to two main factors: (i) the shortcomings of CAC datasets, which primarily consist of images containing objects from a single class, and (ii) the limitations of current counting performance evaluators, which are based on traditional class-specific counting and focus solely on counting errors. To fill this gap, we introduce the Prompt-Aware Counting (PrACo) benchmark. It comprises two targeted tests coupled with evaluation metrics specifically designed to quantitatively measure the robustness and trustworthiness of existing prompt-based CAC models. We evaluate state-of-the-art methods and demonstrate that, although some achieve impressive results on standard class-specific counting metrics, they exhibit a significant deficiency in understanding the input prompt, indicating the need for more careful training procedures or revised designs. The code for reproducing our results is available at https://github.com/ciampluca/PrACo.
**footnotetext: Corresponding authors, they contributed equally to this work.
Figure 1: Prompt-based counting models – CounTX [2] in this example – exhibit difficulties in accurately interpreting user-provided texts that specify object classes to be counted. The confusion occurs even between classes that are semantically very distinct – like marbles and elephants. In some cases, the count of classes not present in the image is even higher than that for the ground-truth object category (highlighted in orange).
@@ 1 Introduction
Class-agnostic counting (CAC) aims to count instances of arbitrary objects beyond the set of categories seen at training [25]. This emerging trend overcomes the limitations affecting long-standing class-specific counting methods, which require individually trained networks for each object type – e.g., people [26, 28, 20, 4, 13, 10], vehicles [33, 1, 7], or cells [9, 6, 5, 22]. In contrast, CAC draws inspiration from humans’ instinctive ability to discern what merits counting when confronted with unfamiliar objects, and it enables users to specify arbitrary categories of interest at inference time without needing model retraining or new annotated data.
@@
Target object classes in CAC can be specified by users with visual exemplars, i.e., bounding boxes surrounding object samples in the image [29, 31, 19, 11, 32], or with text prompts, i.e., textual object descriptions [30, 2, 14, 15, 27, 23, 18]. Techniques leveraging visual exemplars currently outperform prompt-based methods, as exemplars offer more detailed information than text, providing inherent information about the visual appearance of the objects. Nevertheless, prompt-based methods enhance flexibility and improve the capabilities and generality of counting models requiring reduced human intervention – humans do not have to specify bounding boxes as input. These methods leverage the models’ natural language understanding and work hand-in-hand with popular vision-language foundation models such as CLIP [24], thus representing highly appealing solutions [2, 23].
@@
However, in this paper, we figured out that current benchmarks exploited for assessing prompt-based CAC approaches are affected by severe limitations hampering their proper evaluation and the development of new, more performing solutions. As shown in Fig. 1, we empirically verified that some state-of-the-art prompt-based CAC approaches are not always able to truly understand which object class has to be counted from the textual description. In contrast, they tend to count object instances belonging to the predominant class despite the prompt meaning. This is a critical failure in real-world scenarios if we imagine that such systems are usually deployed to count specific objects – e.g., imagine a surveillance scenario in which the officer wants to count the number of pedestrians transiting on the street, but at that moment, only cars are present. We argue that this lack relies on two key factors: (i) the shortcomings affecting the main CAC datasets and (ii) the limitations affecting current counting performance evaluators. Specifically, most of the samples making up the CAC datasets contain objects belonging to a single class, making it hard to evaluate the model’s robustness to discriminate different object classes within an image. Furthermore, the metrics exploited for measuring the performance of CAC models are inherited from class-specific counting and do not consider some key factors of prompt-based CAC, focusing only on the counting error and disregarding assessing the model’s trustworthiness to understand the object category described by the textual prompt.
@@
In this work, we propose Prompt-Aware Counting (PrACo), a novel benchmark designed to quantitatively evaluate the robustness and trustworthiness of prompt-based CAC approaches, tackling the two above-mentioned limitations. The benchmark includes two tailored tests coupled with ad-hoc evaluation metrics: (i) a negative-label test, which probes single-class images by using prompts that refer to absent classes, and (ii) a mosaic test, which evaluates artificially created mosaicked images built by stitching together pairs of single-class images, where one object category serves as a distractor to the category described by the textual prompt. To validate our new benchmark, we assess several recent SOTA prompt-based CAC techniques, and we show that some of them show a notable weakness in understanding the object class textual descriptions despite achieving top performance on standard class-specific metrics. This highlights the necessity for more refined training procedures or reconsideration of their architectural designs.
@@
  • We empirically figured out that the evaluation of state-of-the-art prompt-based class-agnostic counting methods is limited by current datasets and metrics, hindering the development of more effective solutions.
  • We propose PrACo, a novel benchmark that introduces a set of ad-hoc evaluation protocols and associated metrics to evaluate the robustness and trustworthiness of existing CAC models to count objects belonging to classes described with natural language.
  • We exploit PrACo to show that many recent state-of-the-art methods exhibit a remarkable deficiency in understanding objects to be counted, although achieving top results in standard class-specific counting metrics.
@@ 2.1 Class-specific Object Counting
Object counting is one of the core tasks in computer vision due to its widespread applicability to many real-world applications. Thus, many methods have been proposed to count specific object categories, such as people [26, 28, 20, 4, 13], cars [33, 1, 7], insects [3, 8], or cells [9, 6]. Approaches that regress and sum density maps have emerged to be more accurate in crowded scenarios [26, 28, 20, 4, 13, 3] than techniques relying on prior detection of object instances [1, 7]. Either way, the main drawback of these methods lies in requiring individually trained networks and, consequently, labeled datasets for each object type.
@@ 2.2 Prompt-based Class-agnostic Counting
Recently, class-agnostic counting (CAC) has generalized object counting to open-world settings, where users can specify arbitrary object classes never seen during model training by exploiting (i) visual exemplars, i.e., bounding boxes around the objects of interest within the same input image [29, 31, 19, 11, 32], or (ii) textual prompts, i.e., natural-language descriptions of the object class [30, 2, 14, 15, 27, 23, 18]. The first setting provides the best performance: methods such as CACViT [29], SSD [31], CounTR [19], LOCA [11], and SAFECount [32] reach state-of-the-art in all CAC benchmarks. In contrast, prompt-based CAC approaches are particularly appealing since they enhance overall flexibility even if they achieve lower performance.
@@
Most of the SOTA prompt-based CAC approaches rely on vision-language foundation models that map text-image pairs to a joint embedding space. The produced features are then passed to a decoder in charge of producing density maps. The authors in [30] trained a conditional variational autoencoder (VAE) based on the popular CLIP model [24] to generate visual exemplar prototypes conditioned with their semantic embedding encoded in the provided category name. Concurrently, [2] proposed CounTX that instead is trained end-to-end and produces object counts directly, skipping the intermediate visual exemplar proposal step. Being based on CLIP to measure the similarity between image patches and class descriptions, it also accepts a more detailed specification of the target object to count (rather than simply using a class name). Similarly, CLIP-Count [14] relies on CLIP to align the text embedding with dense visual features. However, the authors also designed a hierarchical patch-text interaction module to propagate semantic information across different resolution levels of visual features. Another similar CLIP-based model trained end-to-end is VLCounter [15]. Here, the authors incorporated three modules to efficiently finetune CLIP for the counting task and to exploit intermediate features across different encoding layers of CLIP in the decoding stage. Very recently, DAVE [23] proposed a two-stage detect-and-verify paradigm. In the first stage, they estimate a high-recall set of candidate detections, which are then analyzed and filtered in the second verification step, relying on unsupervised clustering and CLIP embedding. Differently from the above-described architectures, TFPOC [27] employed a detection-based technique, leveraging the popular SAM [16] for instance segmentation. The authors proposed a two-stage approach. In the first stage, they exploited CLIP-Surgery [17], an enhanced version of CLIP, to produce visual exemplar prototypes by computing the similarity between images and text representations. In the second step, they computed a similarity map between image features and the masks corresponding to the reference objects produced by SAM prompted with the bounding boxes from the first stage.
All of these prompt-based methods derive from the literature on exemplar-based CAC, where elements to be counted are chosen by selecting some exemplars from the given image. As a result, prompt-based CAC methods also assume the object is present in the image. Our goal is to test a scenario where this assumption is relaxed, allowing for the possibility that (i) the object may not be present in the image, (ii) the image may contain more than one object class, or even (iii) the user provides a misleading or ambiguous query. We empirically demonstrate that prompt-based CAC models are activated by non-present categories and, therefore, are unable to truly understand object categories to be counted from textual descriptions.
@@ 2.3 Dataset and Metrics
The gold standard dataset for CAC is FSC-147 [25]. It contains 6,135 images across 147 object categories, with 89 categories used for training, 29 for validation, and 29 for testing, with no class overlap between these subsets. Annotations consist of dots over the approximate centroids of each object instance, as usual for the counting task – ground truth density maps are created by superimposing Gaussian kernels centered on these dots. Furthermore, the authors provided three bounding boxes per image that localize the exemplars and the natural language name of the object category to which they belong – the category is unique for each image, and it has been chosen arbitrarily if multiple categories were present. However, cases in which multiple object categories are present in the same image are just a few, and this represents an inherent limitation of FSC-147, making it hard to evaluate the model’s robustness to discriminate different object classes within an image. To address this shortcoming, two other datasets have been proposed: OmniCount-191 [21] and MCAC [12]. These datasets include multiple object categories within the same images. However, the first one is not publicly available, while MCAC lacks suitable annotations for prompt-based approaches.
Standard counting metrics in the literature are the mean absolute error (MAE) and the root mean squared error (RMSE), defined as $\text{MAE}=\frac{1}{N}\sum_{n=1}^{N}\left|\tilde{c}_{n}-c_{n}\right|$ and $\text{RMSE}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(\tilde{c}_{n}-c_{n})^{2}}$, where $N$ is the number of test images, and $\tilde{c}_{n}$ and $c_{n}$ are the ground truth and predicted counts, respectively. Another performance evaluator, used less frequently, is represented by the mean absolute percentage error (MAPE), which is essentially a normalized MAE and is defined as $\text{MAPE}=\frac{1}{N}\sum_{n=1}^{N}\frac{|\tilde{c}_{n}-c_{n}|}{\tilde{c}_{n}}$. However, these evaluators were inherited from the class-specific counting task and exhibit severe limitations particularly evident in prompt-based CAC. Specifically, they do not take into account which object category has to be counted nor the model’s ability to understand the provided textual prompt.
@@ 3.1 Overview
We assume to have a collection of $N$ tuples $\mathcal{X}=\{(I_{1},P_{1}),(I_{2},P_{2}),\dots,(I_{N},P_{N})\}$, where $I_{i}$ is an image and $P_{i}$ is the class name of the objects within $I_{i}$, i.e., the description of the object class present in it. Thus, each image contains objects belonging to only a single class.
We define $P_{i}$ as the positive class and all the $\{P_{j}\}_{j\neq i}$ as the negative classes for the given image $I_{i}$. Without loss of generality, we also assume that each class is represented by only one image in the dataset, even though, in real-world scenarios, datasets typically contain multiple images of the same object class. Finally, we assume to have a prompt-based CAC model $\mathcal{M}(I,P)$, which takes as input a generic image $I$ together with an arbitrary prompt $P$ and outputs an estimated count $c$ of the object instances belonging to the object class described by $P$. In this setting, PrACo introduces two test suites, which are detailed in the following sections: (i) a negative-label test and (ii) a mosaic test.
@@ 3.2 Negative-label test
The negative-label test evaluates single-class images by prompting the model with textual descriptions referring to absent object classes. Formally, for all the elements $\{I_{i},P_{j}\}_{i,j=1}^{N}$, we compute $c_{ij}=\mathcal{M}(I_{i},P_{j})$ to obtain the predicted count for each image when prompted with all the available class descriptions in the dataset. The optimal test outcome can be formalized as follows:
@@
where $\tilde{c}_{i}$ is the ground truth count for the image $I_{i}$.
The ideal situation is illustrated in the left part of Fig. 2. Intuitively, standard class-specific counting metrics consider only the count predictions concerning the positive class – the orange diagonal in the matrix of Fig. 2. Conversely, we account for the count predictions concerning the negative classes – the remaining non-diagonal elements in the matrix of Fig. 2. To quantitatively assess this test, we introduce two ad-hoc metrics.
@@
NMN is the absolute counting error computed by prompting the model with the negative classes, normalized by the ground truth of the positive class. It is computed as follows (more details in the supplementary material):
@@
PCCN mixes positive and negative class predictions, providing an overall quantitative assessment of strong failures in the model. Formally, it is defined as follows:
@@
where $d^{\text{neg}}_{i}=|\frac{1}{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}c_{ij}-\tilde{c}_{i}|$ is the absolute difference between the mean of the model’s negative classes predictions and the ground truth for the $i$-th image, $d^{\text{pos}}_{i}=|c_{ii}-\tilde{c_{i}}|$ is the absolute difference between the model’s positive class prediction and the ground truth count for the $i$-th image, and $\mathbb{I}(\cdot)$ is an indicator function, which equals 1 if the condition inside it is true, and 0 otherwise. Thus, PCCN measures the percentage of data samples for which the model produces a positive class count estimate that is closer to the ground truth compared to the mean of the negative class count estimations. Intuitively, a mean of the negative class estimations that is closer to the ground truth than a positive class estimate may indicate that the model is strongly biased toward counting negative classes over the positive class, which in turn may be due to the model completely losing the semantics of the given prompt. It follows that the higher the PCCN, the better the resulting model.